# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v2

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a principled framework for constructing well-calibrated prediction intervals for non-stationary time series. TCP integrates a machine learning-based quantile forecaster with an online conformal calibration layer. This layer’s thresholds are updated via a modified Robbins-Monro scheme, allowing the model to dynamically adapt to volatility clustering and regime shifts without rigid parametric assumptions. We benchmark TCP against GARCH, Historical Simulation, and static Quantile Regression across diverse financial assets. Our empirical results reveal a critical flaw in static methods: while sharp, Quantile Regression is poorly calibrated, systematically over-covering the nominal 95% target. In contrast, TCP’s adaptive mechanism actively works to achieve the correct coverage level, successfully navigating the coverage-sharpness tradeoff. Visualizations during the 2020 market crash confirm TCP’s superior adaptive response, and a comprehensive sensitivity analysis demonstrates the framework’s robustness to hyperparameter choices. Overall, TCP offers a practical and theoretically-grounded solution to the central challenge of calibrated uncertainty quantification for time series under distribution shift, advancing the interface between statistical inference and machine learning.
@@ 1.   Introduction
Financial risk estimation is more than a regulatory checkbox. It’s a foundational tool for ensuring market stability and investor confidence (Markowitz, 1952; J.P. Morgan, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Value-at-Risk (VaR) and Expected Shortfall (ES) methods, for instance, typically assume Gaussian or other parametric return distributions (Hull & White, 1998), which systematically underestimate extreme losses. Events like the 2008 global financial crisis (Brunnermeier, 2009) and the COVID-19 crash (Baker et al., 2020) exposed the fragility of these assumptions, especially in the tails.
@@
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some of this volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
@@
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on the assumption of exchangeability, a condition often violated by time series data with temporal dependencies. Several recent works attempt to address this by extending conformal methods to sequential or temporally dependent settings (Xu & Xie, 2021; Stankeviciute et al., 2021), but a robust and practical solution for financial markets remains an open challenge.
@@
This paper closes that gap. We introduce Temporal Conformal Prediction (TCP), a real-time, adaptive framework for constructing well-calibrated prediction intervals for financial time series. TCP combines a modern machine learning-based quantile forecaster with an online conformal calibration layer governed by a modified Robbins-Monro scheme. Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
@@
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR). Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically over-covers the target 95% confidence level. TCP, by contrast, demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate. This adaptive property, which we visualize during the COVID-19 crash, makes TCP a more reliable and principled framework for real-world risk management. Furthermore, a comprehensive sensitivity analysis underscores the robustness of our framework to key hyperparameter choices.
@@
We organize the paper as follows. Section 2 surveys related work. Section 3 establishes our mathematical notation. Section 4 lays out the TCP theory. Section 5 details our empirical setup, including the model architectures and evaluation framework. Section 6 presents our main findings, including the benchmark comparisons, visualizations, and hyperparameter sensitivity analysis. Finally, Section 7 wraps up with practical takeaways and future directions.
@@ 2.   Related Work
Before detailing our proposed TCP framework, we situate our work within the existing literature. This section reviews three key research areas that inform our approach: traditional financial risk models, the evolution of time series forecasting with machine learning, and the theoretical foundations of conformal prediction.
Traditional Financial Risk Models. Quantitative risk management traces its roots to Markowitz’s portfolio theory (Markowitz, 1952) and was formalized around the Value-at-Risk concept with J.P. Morgan’s RiskMetrics in 1996 (Morgan, 1996). In practice, three main VaR estimation paradigms have dominated: parametric approaches like GARCH-style models, which assume a specific return distribution; non-parametric Historical Simulation, which relies on past returns; and Monte Carlo methods, which simulate future scenarios. Beyond VaR, Conditional Value-at-Risk (CVaR) was proposed as a coherent alternative (Rockafellar & Uryasev, 2000). However, a common weakness is that these traditional approaches often fail backtests during market crises (Kupiec, 1995; Christoffersen, 1998), highlighting their blind spot when tail events dominate.
Advanced Time Series and Machine Learning Models. To better capture the dynamics of financial data, the GARCH family of models became foundational for modeling time-varying variance and volatility clustering (Engle, 1982; Bollerslev, 1986). Extreme Value Theory (EVT) offered a principled way to model the tails of distributions directly (Pickands, 1975; McNeil & Frey, 2000). More recently, machine learning has injected fresh momentum into risk estimation. Quantile regression (Koenker & Bassett, 1978) directly targets conditional quantiles without assuming a distribution, with modern variants including gradient-boosting forests (Meinshausen, 2006). Deep learning, especially LSTM networks, has also shown promise in capturing complex temporal dependencies (Fischer & Krauss, 2018).
Conformal Prediction Theory. Bridging these worlds is conformal prediction (Vovk et al., 2005), a framework that provides distribution-free, finite-sample prediction intervals. However, classical conformal methods assume exchangeability, a condition that breaks down in time series. Recent advances have sought to address this limitation by introducing adaptive calibration to handle distribution shift (Gibbs & Candès, 2021) and by developing extensions specifically for temporal data (Xu & Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations, proposing a practical and robust adaptive conformal framework tailored specifically for the challenges of financial risk forecasting.
@@ 3.   Notation
We denote the price of a financial asset at time $t$ as $P_{t}$, and its corresponding daily log-return as $r_{t}$. The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $[\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate. The feature space for our models includes lagged returns and realized volatility $\sigma_{t}$.
Learning and Prediction: For a general learning problem with $n$ observations, we consider feature-label pairs as $(X_{i},Y_{i})$ and a generic point prediction model as $\widehat{f}(\cdot)$. Our time series evaluation runs for $T$ total time steps. We use $\mathcal{F}_{t}$ to represent the filtration (information set) available up to time $t$. The true, unknown $\tau$-quantile of the return distribution is $q_{\tau}$. Our quantile regression model, $\mathcal{F}_{\tau}$, produces a data-driven estimate of this quantile, denoted $\widehat{q}_{\tau}$.
Conformal Prediction: Our conformal prediction framework relies on a non-conformity score function, $A(\cdot,\cdot)$, which produces a score $\alpha_{i}$ for each observation. For our method, we use the model residuals as scores, $\varepsilon_{i}$. This process generates a conformal prediction set, $\Gamma_{\alpha}$. The method’s local adaptation is managed using a rolling window of size $w$, which informs the adaptive conformal threshold at time $t$, denoted $C_{t}$.
@@
[…]
@@ 4.1 Problem Formulation
We begin by framing the central task of this paper. We consider a univariate financial time series of daily log-returns $\{r_{1},r_{2},\dots,r_{t}\}$, where $r_{t}=\log(P_{t}/P_{t-1})$ and $P_{t}$ is the asset price at time $t$. Our objective is to construct a prediction interval $[\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$. Traditional methods often fail to achieve this in the presence of non-stationarity, motivating our distribution-free approach.
@@ 4.2 Classical Conformal Prediction
A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model. A full proof is provided in Appendix A.
@@ 4.3 Temporal Conformal Prediction (TCP)
Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$
@@
1:Time series of returns $\{r_{1},\dots,r_{t}\}$, window size $w$, miscoverage rate $\alpha$.
2:Train Quantile Models: Fit quantile regression models $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ on the training set $\{r_{i-w},\dots,r_{i-1}\}$ for $i=t-w+1,\dots,t$.
3:Compute Non-Conformity Scores: For each $i$ in the window, compute the scores (residuals) $\varepsilon_{i,\alpha/2}=r_{i}-\widehat{q}_{\alpha/2}(X_{i})$ and $\varepsilon_{i,1-\alpha/2}=\widehat{q}_{1-\alpha/2}(X_{i})-r_{i}$.
4:Calculate Conformal Threshold: Find the $(1-\alpha)$-quantile of the non-conformity scores from the window: $C_{t}=\text{quantile}(\{\varepsilon_{i,\tau}\}_{i=t-w+1}^{t},1-\alpha)$.
5:Form Prediction Interval: Predict the quantiles for the next step, $\widehat{q}_{\alpha/2}(X_{t+1})$ and $\widehat{q}_{1-\alpha/2}(X_{t+1})$. Construct the interval:
$[\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\quad\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$
6:Prediction interval $[\ell_{t+1},u_{t+1}]$.
4.4 Adaptive Calibration
While the window-based approach in Algorithm 1 provides local adaptation, systematic drifts can still lead to miscalibration. To address this, we introduce an online learning mechanism that actively adjusts the conformal threshold to maintain the target coverage rate over time.
Definition 4.2 (Coverage Error)
At each time $t$, the coverage error is defined as $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$.
We then adjust the threshold using a Robbins-Monro style update:
$C_{t+1}=C_{t}+\gamma_{t}e_{t},\quad\text{where}\quad\gamma_{t}=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}}$
is a decaying learning rate with hyperparameters $\gamma_{0}>0,\lambda>0,\beta\in(0.5,1]$, ensuring that $\sum_{t}\gamma_{t}=\infty$ and $\sum_{t}\gamma_{t}^{2}<\infty$ (Robbins & Monro, 1951; Kushner & Yin, 2003).
Theorem 4.2 (Asymptotic Coverage)
Under the update rule in Equation (5) and standard ergodicity assumptions, the long-run average coverage converges to the nominal level:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha\quad a.s.$
The proof is detailed in Appendix A.
Clarification on the Adaptive Update Mechanism
Our online update rule for the conformal thresholds, $C_{t}$, is a practical modification of the classic Robbins-Monro (Robbins & Monro, 1951) stochastic approximation scheme. The standard algorithm adjusts the thresholds only in the event of a coverage error. While this guarantees asymptotic convergence, it can be slow to adapt in practice. Specifically, during a volatile period, the intervals will widen but may fail to narrow again once the market stabilizes, as the absence of coverage errors provides no signal to reduce the thresholds.
To address this and improve the crucial coverage-sharpness tradeoff, we introduce a small decay heuristic. When an observation $r_{t}$ falls correctly within the prediction interval, we symmetrically reduce the conformal thresholds by a small fraction of the learning rate. This modification acts as a “forgetting factor,” preventing the intervals from becoming permanently wide and allowing them to become sharper during periods of sustained low volatility. While this practical enhancement is a deviation from the classic formulation and is not covered by our formal proof, our empirical results and visualizations demonstrate its effectiveness in creating more responsive and efficient prediction intervals in dynamic financial markets.

### v3

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, rolling QR yields the sharpest intervals but is materially under‐calibrated (e.g., S&P 500: 86.3% vs. 95% target). Second, TCP (and TCP–RM) achieves near‐nominal coverage while being substantially narrower than Historical Simulation (e.g., S&P 500: $29\%$ width). Third, the RM update improves calibration with negligible width cost. Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step‐size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 1.  INTRODUCTION
Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P. Morgan/Reuters, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk. Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000). Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
@@
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
@@
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies. Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
@@
This paper closes that gap. We introduce Temporal Conformal Prediction (TCP), a real-time, adaptive framework for constructing well-calibrated prediction intervals for financial time series. TCP combines a modern machine-learning quantile forecaster with an online conformal calibration layer governed by a modified Robbins–Monro scheme. Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
@@
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR). Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically misses the 95% target. By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate. This adaptive property, which we visualize during the March 2020 crash, makes TCP a more reliable and principled framework for real-world risk management. A comprehensive sensitivity analysis further underscores the robustness of our framework to key hyperparameters.
@@
We organize the paper as follows. Section 2 surveys related work. Section 3 establishes our mathematical notation. Section 4 lays out the TCP theory. Section 5 details our empirical setup, including model architectures and the evaluation framework. Section 6 presents our main findings, including benchmark comparisons, visualizations, and hyperparameter sensitivity analysis. Finally, Section 7 offers practical takeaways and future directions.
@@ 2.  RELATED WORK
Before detailing our proposed TCP framework, we situate our work within the existing literature. We review three areas that inform our approach: traditional financial risk models, the evolution of time-series forecasting with machine learning, and the theoretical foundations of conformal prediction.
Traditional Financial Risk Models. Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P. Morgan’s RiskMetrics (J.P. Morgan/Reuters, 1996). In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004). Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000). However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
Advanced Time-Series and Machine-Learning Models. To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993). For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (III, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000). On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018). We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
Conformal Prediction Theory. Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005). Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
@@ 3.  NOTATION
We denote the price of a financial asset at time $t$ as $P_{t}$, and its corresponding daily log-return as $r_{t}$. The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate. The feature space for our models includes lagged returns and realized volatility $\sigma_{t}$.
[…]
Conformal Prediction (two-sided, split). At each time $t$ we use a rolling window of size $w$ split into a training slice $\mathcal{T}_{t}$ and a calibration slice $\mathcal{C}_{t}$ with $|\mathcal{T}_{t}|+|\mathcal{C}_{t}|=w$. Quantile forecasters $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only (no leakage). Two-sided nonconformity scores on the calibration slice are
$s_{i}=\max\{\widehat{q}_{\alpha/2}(X_{i})-r_{i},\;r_{i}-\widehat{q}_{1-\alpha/2}(X_{i}),\;0\},\quad i\in\mathcal{C}_{t}.$
Let $m=|\mathcal{C}_{t}|$ and $s_{(1)}\leq\cdots\leq s_{(m)}$ be the order statistics. We set the conformal threshold $C_{t}=s_{(k)}$ with $k=\lceil(m+1)(1-\alpha)\rceil$, and the next-step interval $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\;\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$. For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice $|\mathcal{T}_{t}|{=}w_{\mathrm{tr}}$ (no calibration slice), i.e., it is rolling and out-of-sample.
@@
Mathematical Operators: Finally, we use standard mathematical operators, including $\Pr(\cdot)$ for probability, $\mathbb{E}[\cdot]$ for expectation, $\mathbf{1}(\cdot)$ for the indicator function, $\mathrm{sign}(\cdot)$ for the sign function, and $\lfloor\cdot\rfloor$ for the floor function.
Variants used (terminology). We consider two closely related procedures. TCP denotes split–conformal prediction on a rolling window: the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ (size $w_{\mathrm{tr}}$) and a calibration slice $\mathcal{C}_{t}$ (size $w_{\mathrm{cal}}$), $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$. Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only and the two-sided split–conformal threshold $C_{t}$ is computed from $\mathcal{C}_{t}$, yielding $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$. TCP-RM augments TCP with an online Robbins–Monro (RM) offset $C_{t}^{\mathrm{RM}}$ updated by $C_{t+1}^{\mathrm{RM}}=C_{t}^{\mathrm{RM}}+\gamma_{t}\{\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha\}$, with $\gamma_{t}=\gamma_{0}(1+\lambda t)^{-\beta}$. The effective threshold is $C_{t}^{\mathrm{eff}}=C_{t}+C_{t}^{\mathrm{RM}}$ and the interval becomes $\widehat{q}_{\alpha/2}(X_{t+1})-C_{t}^{\mathrm{eff}},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}^{\mathrm{eff}}]$. For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice ($|\mathcal{T}_{t}|=w_{\mathrm{tr}}$) and no calibration slice (rolling, out-of-sample).
@@ 4.1 Problem Formulation
We begin by framing the central task of this paper. We consider a univariate financial time series of daily log-returns $\{r_{1},r_{2},\dots,r_{t}\}$, where $r_{t}=\log(P_{t}/P_{t-1})$ and $P_{t}$ is the asset price at time $t$. Our objective is to construct a prediction interval $\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$. Traditional methods often fail to achieve this in the presence of non-stationarity, motivating our distribution-free approach.
@@ 4.2 Classical Conformal Prediction
A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model.
A full proof is provided in Appendix A.
@@ 4.3 Temporal Conformal Prediction (TCP)
Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$ (split-conformal, two-sided)
@@
0: Returns $\{r_{1},\dots,r_{t}\}$; features $\{X_{1},\dots,X_{t}\}$; window $w$ with $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$; miscoverage $\alpha\in(0,1)$
1:Define slices: $\mathcal{W}_{t}=\{t-w+1,\dots,t\}$; $\mathcal{T}_{t}=\{t-w+1,\dots,t-w_{\mathrm{cal}}\}$; $\mathcal{C}_{t}=\{t-w_{\mathrm{cal}}+1,\dots,t\}$
2:Fit forecasters on $\mathcal{T}_{t}$: train $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ on $\{(X_{i},r_{i}):i\in\mathcal{T}_{t}\}$
3:Compute two-sided nonconformity on $\mathcal{C}_{t}$:
4:for each $i\in\mathcal{C}_{t}$ do
5:  $s_{i}\leftarrow\max\{\widehat{q}_{\alpha/2}(X_{i})-r_{i},\;r_{i}-\widehat{q}_{1-\alpha/2}(X_{i}),\;0\}$
[…]
7:Conformal threshold: $k\leftarrow\lceil(m+1)(1-\alpha)\rceil$;  $C_{t}\leftarrow s_{(k)}$
8:Form next-step interval:
[…]
9: Prediction interval $\ell_{t+1},u_{t+1}]$
4.4 Adaptive Calibration (TCP vs. TCP-RM)
We consider two variants. TCP is split-conformal only: at each time $t$, the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ and a calibration slice $\mathcal{C}_{t}$ with $|\mathcal{T}_{t}|+|\mathcal{C}_{t}|=w$. Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ and the two-sided nonconformity scores on $\mathcal{C}_{t}$ yield the split-conformal threshold $C_{t}=s_{(k)}$, $k=\lceil(|\mathcal{C}_{t}|+1)(1-\alpha)\rceil$. TCP-RM augments TCP with a single online offset $C^{\mathrm{RM}}_{t}$ updated by Robbins–Monro:
$C^{\mathrm{RM}}_{t+1}$ $=C^{\mathrm{RM}}_{t}+\gamma_{t}\Big(\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha\Big),$ (1)
$\gamma_{t}$ $=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}},\ \beta\in(1/2,1].$ (2)
The effective threshold is $C^{\mathrm{eff}}_{t}=C_{t}+C^{\mathrm{RM}}_{t}$, giving $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C^{\mathrm{eff}}_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C^{\mathrm{eff}}_{t}]$. No manual decay or heuristic shrinkage is used.
Assumptions. (A1) (Bounded scores) $\{s_{i}\}$ are a.s. bounded. (A2) (Local mixing) The process is $\beta$-mixing so that $e_{t}-\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]$ is a martingale-difference with bounded variance. (A3) (Monotone coverage in $C$) $g(C)=\Pr(r_{t}\in[\ell_{t},u_{t}]\mid C)$ is continuous and strictly increasing near the unique $C^{\star}$ with $g(C^{\star})=1-\alpha$. (A4) (Stepsizes) $\gamma_{t}>0$, $\sum_{t}\gamma_{t}=\infty$, $\sum_{t}\gamma_{t}^{2}<\infty$.
Under (A1)–(A4), TCP provides local finite-sample validity on $\mathcal{C}_{t}$; TCP-RM adds asymptotic calibration in time via (1).
4.5 Proof of Asymptotic Calibration
Theorem 4.2
Under the Robbins–Monro update (1) and (A1)–(A4), $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s., and $C_{t}\to C^{\star}$ a.s.
Full proof is given in Appendix A.
Remark 4.1 (Practical stability)
To enforce (A1) and keep the iterates bounded, we only enforce nonnegativity of the effective threshold by projecting the RM offset:
$C^{\mathrm{RM}}_{t+1}\leftarrow\max\Big\{\,C^{\mathrm{RM}}_{t}+\gamma_{t}\big(\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha\big),\ -\,C_{t}\Big\},$
so that $C^{\mathrm{eff}}_{t+1}=C_{t}+C^{\mathrm{RM}}_{t+1}\geq 0$. An additional upper cap $S_{t}$ (e.g., a high empirical quantile of the calibration scores $s_{i}$) can be included, but we did not use it in our experiments. This projection is standard in stochastic approximation and does not change the prediction rule.